Oak View Chemicals produces two main products, M-10 and M-20, and one by-product, B-60, from a single input, Chem-X. Products M-10 and M-20 can either be sold at split-off or processed further and sold. A given batch begins with 1,220 gallons of Chem-X with a cost of $109,800. Additional information regarding a batch follows.

If Processed Further
Product Units Produced Unit Sales Value at Split-off Additional Costs (Per Unit) Sales Value (Per Unit)
M-10 26,840 $ 5.00 $ 2.00 $ 6.00
M-20 17,080 $ 10.00 $ 4.00 $ 16.00
B-60 6,100 $ 0.50 NA NA

Oak View Chemicals uses the net realizable at split-off approach to allocate joint costs and treats the sales value of the by-product B-60 as other income.


Required:

The joint cost allocated to product M-10 at Oak View Chemicals for a given batch is:

Multiple Choice

  • $67,100

  • $48,312

  • $46,970

  • $40,713

  • None of the above.
Answer #1

(B) $48,312

Total Joint Cost = $109800

Sale value at Split off Point:

M-10 = 26840×$5= $134200

M-20 = 17080×$10= $170800

Total sales value at split off point = $134200+$170800=$305000

Joint cost allocated to M-10 = Total Joint cost * Sale value of M-10 at split off / Total sales value at split off

=$109800* $305000/*134200

=$48312

(As the sales value of By-Product B-60 is treated as other income, it is not reduced here from total Joint cost.)